Credenza Industries is expected to pay a dividend of $1.15 at the end of the coming year. It is expected to sell for $68 at the end of the year. If its equity cost of capital is 88​%, what is the expected capital gain from the sale of this stock at the end of the coming​ year?

A. $64.03

B. $62.96

C. $5.04

D. $ 3.97
Answer #1

Answer: Option D is correct
Present value=Present value of the dividend Dividend and the selling price at cost of capital of 8%
=1.15/(1+8%)+68/(1+8%)
=64.02777778

Capital gain=Selling price-Present value
=68-64.02777778=3.97222222 or $3.97 (Rounded to two decimal places)